OpenStudy (anonymous):

second derivative of f(x)=-7x*e^(-3.5*x^2)

OpenStudy (anonymous):

Did you find the first derivative yet?

OpenStudy (anonymous):

that is the first derivative

OpenStudy (anonymous):

should I just do product rule? or chain?

OpenStudy (anonymous):

it requires use of the product rule and the chain rule.

OpenStudy (anonymous):

Oh, then you just need to take the derivative once. It is a product, so you will need the product rule. It is also a function composition so you'll need the chain rule also..

OpenStudy (anonymous):

ok thank you that's where I got stumped

OpenStudy (anonymous):

If the function is a product you will need the product rule.

OpenStudy (anonymous):

But if it's a function in a function you need the chain rule, and if it's both, you need both.

OpenStudy (anonymous):

\[f'(x)=-7*e ^{7/2* x ^{2}}-7x*(-7x)*e ^{-7/2 *x ^{2}}=...\] simplify. can you find the second derivative?

OpenStudy (anonymous):

what you wrote is the second derivative that I got but I factored out a -7.

OpenStudy (anonymous):

derivative of e^(-7/2 x^2)=-7/2 *2x * e^(...)=-7x*e^(...)

OpenStudy (anonymous):

are you Ok now?

OpenStudy (anonymous):

yes I am asked to then find where it is concave up or down, so I am setting that second derivative to zero.

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

it will give you inflection points

OpenStudy (anonymous):

i got inflection points: \[x= \pm 1/\sqrt{7}\]